# SPHERE: A Hierarchical Evaluation on Spatial Perception and Reasoning for Vision-Language Models

### v1

@@ 4.1 Setup
In this section, we describe the experimental setup used to evaluate the performance of state-of-the-art open-source VLMs on SPHERE tasks. All experiments were performed on a NVIDIA GeForce RTX 3090 GPU.
Models: Our evaluation encompasses two model categories: general-purpose VLMs and VLMs with enhanced spatial capabilities. The general-purpose VLMs evaluated are Phi-3.5-Vision (4.1B) Abdin et al. (2024), LLaVA-NeXT (7.1B) Liu et al. (2024), InstructBLIP (7.9B) Dai et al. (2023), Idefics2 (8.4B) Hugging Face (2024), Qwen-VL (9.7B) Bai et al. (2023). The VLMs specifically trained for spatial understanding and reasoning are SpatialBot (2.8B) (Cai et al., 2024) and SpaceMantis (8.4B) remyx.ai (2024b). Besides the default version of SpatialBot that accepts both RGB and depth map as inputs, we also include the RGB-only version of SpatialBot as SpatialBot-RGB. SpaceMantis remyx.ai (2024b) is fine-tuned from Mantis-8B-siglip-llama3 remyx.ai (2024b) following techniques described in (Chen et al., 2024). All model implementations follow their released versions on Hugging Face.
@@ 4.2 Results
Table 2 shows model accuracy on each of the SPHERE tasks. The results highlight significant room for improvement, indicating that the current models struggle to achieve high performance in these tasks. Interestingly, we observe that the smaller models, Phi-3.5-Vision (4.1B) and SpatialBot (2.8B), outperform larger models in terms of overall accuracy.
For the single-skill tasks, SpatialBot has the highest accuracy (62.5%), followed closely by Phi-3.5-Vision (62.0%). Idefics2 has the lowest accuracy (47.1%) and performs worse than the baseline on Position and Distance. Out of the 4 tasks, most models perform the worst on Distance, indicating that current VLMs still have difficulty understanding distance and proximity.
For the multi-skill tasks, Phi-3.5-Vision has the highest accuracy (39.9%) and Idefics2 has the lowest accuracy (30.2%). The multi-skill tasks are markedly more challenging than the single-skill tasks. In particular, we observe that all models except for SpaceMantis perform worse than baseline random guessing on Distance + Size. The average accuracy on Size is 63.5% while the average accuracy on Distance + Size is 36.3%. This result reflects the reliance of the VLMs on the 2D representation of objects in an image, and measures object size as being directly proportional to the number of pixels occupied. The failure to incorporate the principle of size constancy may cause VLMs to struggle in 3D spaces.
For the reasoning tasks, Phi-3.5-Vision and SpatialBot have the highest accuracy (56.8%) and Qwen-VL has the lowest accuracy (48.7%). Most models perform at the level of random guessing, suggesting that their training datasets may lack sufficient examples emphasizing reasoning skills.
We refer readers to Table 6 in the Appendix B for detailed validity scores of the models. The validity scores reflect the instruction-following capabilities of the VLMs. Most models effectively follow instructions and provide valid responses, such as numerical answers for counting-related questions and selecting an option from the provided list for multiple-choice questions. Idefics2 produces the highest proportion of invalid responses, approximately 8.7%, across all tasks. LLaVA-NeXT and Qwen-VL show a high proportion of invalid responses for certain tasks. Common types of invalid responses include providing an image caption instead of addressing the question, refusing to answer, or offering an answer that is not among the given options.
@@ 4.3 Further Analysis
Bias toward allocentric or egocentric viewpoint. We further analyze model performance on position-related tasks by distinguishing between allocentric and egocentric viewpoints. This analysis provides insights into whether VLMs can comprehend object orientations and reason from both relative (object-to-object) and absolute (viewer-to-object) perspectives. Example questions are shown in Figure 4. From the results in Table 3, we observe that the VLMs are divided in their biases for allocentric or egocentric viewpoint. InstructBLIP, Idefics2 and SpaceMantis show a slight bias toward the egocentric viewpoint. The other models tested demonstrate a significant bias toward the allocentric viewpoint, with performance gaps as large as for Qwen-VL in the Position task, and for LLaVA-NeXT in the Position + Counting task.
@@
From Table 4, we observe that incorporating additional information improves final question accuracy across all models for the object occlusion task, but yields mixed results for the object manipulation task. While the additional information can introduce relevant facts that the models might otherwise overlook, the models sometimes struggle to establish a clear connection between these facts and the specific reasoning they need to conduct. Even with the provided facts, none of the VLMs tested achieved an accuracy above 60%, highlighting a significant deficiency in their reasoning capabilities.
Skill composition. The different performance of the evaluated models across tasks reflect that they have different compositions of skills, and that proficiency in one task does not necessarily translate to proficiency in related tasks. For instance, relatively weaker basic skills do not necessarily imply correspondingly weaker higher-order skills. In Table 2, SpaceMantis has 8.4% lower average accuracy than SpatialBot on the single-skill tasks, but has 5% higher average accuracy than SpatialBot on the multi-skill tasks, and is one of the higher-performing models on the reasoning tasks. In Table 4, on the object manipulation task, Idefics2 and Qwen-VL have attained higher accuracy on the final reasoning questions (53.6% and 49.0%) than on the supposedly easier intermediate perception or understanding questions (47.6% and 48.0%).
@@ 6 Limitations
As we seek to ensure the quality of the SPHERE dataset through manual annotations, the dataset may be limited in scope. The framework and the curated tasks may not encompass all possible spatial understanding and reasoning challenges encountered in practice. In addition, the dataset focuses on static images, which may not fully capture the challenges of dynamic spatial understanding and reasoning in real-world scenarios.
@@ Appendix A Detailed Dataset Statistics
Counting | 5.7879 | 7 | 0.56472 |
Distance + Counting | 9.7151 | 6 | 0.13717 |
Position + Counting | 8.6029 | 5 | 0.12599 |
@@ Appendix B Detailed Evaluation Results
[…]
Table 6 shows the validity scores of the models across all tasks. These scores reflect the instruction-following capabilities of the VLMs. Most models effectively follow instructions and provide valid responses, such as numerical answers for counting-related questions and selecting an option from the provided list for multiple-choice questions. Idefics2 produces the highest proportion of invalid responses, approximately 8.7%, across all tasks. LLaVA-NeXT and Qwen-VL show a high proportion of invalid responses for certain tasks. Common types of invalid responses include providing an image caption instead of addressing the question, refusing to answer, or offering an answer that is not among the given options.

### v2

@@ 4.1 Setup
We describe the experimental setup used to evaluate the performance of state-of-the-art VLMs on SPHERE tasks. All experiments were performed on a NVIDIA GeForce RTX 3090 GPU.
Models: Our evaluation encompasses two model categories: general-purpose VLMs and VLMs with enhanced spatial capabilities. The general-purpose VLMs evaluated are Phi-3.5-Vision Abdin et al. (2024), LLaVA-NeXT Liu et al. (2024), LLaVA-OneVision Li et al. (2024a), InstructBLIP Dai et al. (2023), Idefics2 HuggingFaceM4 (2024), InternVL2.5 Chen et al. (2024b), Qwen-VL Bai et al. (2023), Qwen2-VL Wang et al. (2024), Llama-3.2-Vision Meta (2024), Gemini 2.0 Flash Google (2024), and GPT-4o OpenAI (2024). The VLMs specifically trained for spatial understanding and reasoning are SpatialBot (Cai et al., 2024), SpaceMantis AI (2024b), and SpatialRGPT Cheng et al. (2024). For SpatialRGPT, we use the RGB-only version as the model requires highly accurate mask proposals to make use of depth maps. All model implementations follow their released versions on Hugging Face.
@@ 4.2 Results
Table 3 shows model accuracy on each of the SPHERE tasks. The results highlight significant room for improvement, especially for multi-skill and reasoning tasks.
The proprietary models, GPT-4o and Gemini 2.0 Flash, outperform the open-weights models evaluated in almost all tasks. Interestingly, amongst open-weights models, we observe that the smaller model LLaVA-OneVision (7B) can outperform larger models in terms of overall accuracy, and performs on par with Gemini 2.0 Flash.
For the single-skill tasks, Gemini 2.0 Flash has the highest average accuracy (78.2%). The best performances on Counting and Size are above 80%. The best performances on Position and Distance are between 70% and 80%, indicating that current VLMs still have difficulty in these two skills. In Section 4.3, we analyze how viewpoint affects model performance on understanding positions.
For the multi-skill tasks, GPT-4o has the highest average accuracy (58.6%). The multi-skill tasks are markedly more challenging than the single-skill tasks. In particular, we observe that half of the models perform worse than baseline random guessing on Distance + Size. Qwen2-VL (72B) obtains the highest 69.3% on Distance + Size, and 85.5% on Size. This gap reflects the reliance of the VLMs on the 2D representation of objects in an image, and measures object size as being directly proportional to the number of pixels occupied. The failure to incorporate the principle of size constancy may cause VLMs to struggle in 3D spaces.
For the reasoning tasks, GPT-4o has the highest average accuracy (64.7%). Most models perform close to random guessing, suggesting that their training datasets may lack sufficient examples emphasizing reasoning skills. Notably, Gemini 2.0 Flash has the worst average accuracy (48.6%), below the random guessing baseline.
@@ 4.3 Further Analysis
Bias toward allocentric or egocentric viewpoint. We further analyze model performance on position-related tasks by distinguishing between allocentric and egocentric viewpoints. This analysis provides insights into whether VLMs can comprehend object orientations and reason from both relative (object-to-object) and absolute (viewer-to-object) perspectives. Example questions are shown in Figure 4. From the results in Table 4, we observe that most VLMs demonstrate a significant bias toward the allocentric viewpoint, with performance gaps as large as for Phi-3.5-Vision in the Position task, and for LLaVA-OneVision in the Position + Counting task.
@@
From Table 5, we observe that incorporating additional information improves final question accuracy across all models for the object occlusion task, but yields mixed results for the object manipulation task. Although the additional information can introduce relevant facts that the models might otherwise overlook, the models sometimes struggle to establish a clear connection between these facts and the specific reasoning they need to conduct.
Skill composition. The different performance of the evaluated models across tasks reflect that they have different compositions of skills, and that proficiency in one task does not necessarily translate to proficiency in related tasks. For instance, relatively weaker basic skills do not necessarily imply correspondingly weaker higher-order skills. In Table 3, SpaceMantis has 7.3% lower average accuracy than SpatialBot on the single-skill tasks, but has 1.2% higher average accuracy than SpatialBot on the multi-skill tasks. In Table 5, on the object manipulation task, Idefics2 and Llama-3.2-Vision (11B) have attained higher accuracy on the final reasoning questions (53.3% and 56.1%) than on the supposedly easier intermediate perception or understanding questions (51.1% and 53.7%).
@@ 6 Limitations
As we seek to ensure the quality of the SPHERE dataset through manual annotations, the dataset may be limited in scope. The framework and the curated tasks may not encompass all possible spatial understanding and reasoning challenges encountered in practice. In addition, the dataset focuses on static images, which may not fully capture the challenges of dynamic spatial understanding and reasoning in real-world scenarios. As VLMs advance in spatial perception and reasoning, automatic annotation methods may become more reliable and warrant exploration in future work.
@@ Appendix A Detailed Dataset Statistics
Object concepts.
We assess the diversity of object concepts in the annotations by counting the occurrences of unique nouns. Nouns were extracted from each question using the en_core_web_sm model from spaCy, then manually filtered to exclude propositions and commonly referenced objects in question phrasings, such as ‘camera’, ‘viewer’, ‘photo’, ‘image’. The dataset contains a total of more than 600 object concepts. Figure 10 shows the distribution of object concepts across the dataset for the top 30 concepts.
@@
| Counting | 6.2420 | 7 | 0.51179 |
| Distance + Counting | 7.8268 | 6 | 0.25107 |
| Position + Counting | 7.8879 | 5 | 0.16252 |
@@ Appendix B Detailed Evaluation Results
Table 7 shows the validity scores of the models across all tasks. These scores reflect the instruction-following capabilities of the VLMs. Most models effectively follow instructions and provide valid responses, such as numerical answers for counting-related questions and selecting an option from the provided list for multiple-choice questions. Idefics2 has above 95% validity, and all other models have above 98% validity. Common types of invalid responses include providing an image caption instead of addressing the question, refusing to answer, or offering an answer that is not among the given options.
Model Single-skill Multi-skill Reasoning Pos. Count. Dist. Size Avg P + C D + C D + S Avg Occl. Manip. Avg Overall General purpose VLMs Phi-3.5-Vision (4B) 100.0 100.0 99.7 100.0 99.9 100.0 100.0 99.0 99.7 100.0 100.0 100.0 99.9 LLaVA-NeXT (7B) 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.8 99.6 100.0 100.0 100.0 99.9 LLaVA-OneVision (7B) 99.7 100.0 99.8 98.4 99.5 100.0 100.0 97.3 99.1 99.7 100.0 99.8 99.4 Qwen2-VL (7B) 97.4 100.0 98.3 99.1 98.7 100.0 100.0 98.5 99.5 98.5 99.5 99.0 99.0 InstructBLIP (8B) 97.1 100.0 100.0 99.7 99.2 99.4 100.0 99.9 99.8 99.8 100.0 99.9 99.6 Idefics2 (8B) 92.9 93.0 96.5 91.9 93.6 99.4 96.8 91.4 95.9 98.7 96.0 97.3 95.3 InternVL2.5 (8B) 98.2 100.0 99.4 98.9 99.1 100.0 100.0 99.3 99.8 99.8 100.0 99.9 99.5 Qwen-VL (10B) 99.0 100.0 99.5 100.0 99.6 100.0 100.0 98.4 99.5 96.8 98.8 97.8 99.1 Llama-3.2-Vision (11B) 99.0 100.0 99.0 98.3 99.1 100.0 100.0 98.6 99.5 100.0 99.9 99.9 99.4 Qwen2-VL (72B) 97.6 100.0 99.5 99.0 99.0 100.0 100.0 99.3 99.8 99.6 100.0 99.8 99.5 Llama-3.2-Vision (90B) 95.9 100.0 99.5 98.3 98.4 100.0 100.0 98.3 99.4 100.0 98.8 99.4 99.0 Gemini 2.0 Flash 99.4 100.0 100.0 100.0 99.9 100.0 100.0 98.5 99.5 100.0 100.0 100.0 99.8 GPT-4o 99.2 99.5 98.5 97.0 98.5 100.0 100.0 94.0 98.0 98.5 100.0 99.2 98.5 Spatial VLMs SpatialBot-RGB (3B) 99.2 99.0 99.8 98.4 99.1 100.0 100.0 99.0 99.7 100.0 99.5 99.7 99.4 SpatialBot (3B) 99.2 100.0 100.0 100.0 99.8 100.0 99.4 99.3 99.6 100.0 100.0 100.0 99.8 SpaceMantis (8B) 99.4 100.0 100.0 99.7 99.8 100.0 100.0 98.5 99.5 100.0 100.0 100.0 99.7 SpatialRGPT-RGB (8B) 98.9 100.0 100.0 99.1 99.5 100.0 100.0 99.2 99.7 97.9 95.9 96.9 98.9